OpenStudy (anonymous):

I need help in solving this problem can someone help....The question ask to write the domain and range of the function in interval notation. f(x) sqx-3. Then it ask the range of the function

OpenStudy (anonymous):

\[f(x)=\sqrt{x-3}\]?

OpenStudy (anonymous):

It does give me multipal anwers to choose from

OpenStudy (anonymous):

you cannot take the square root of a negative number, so your first job to find the domain is to solve \(x-3\geq 0\) in one step

OpenStudy (anonymous):

if it is the function i wrote, then one of the answers for the domain should be either \[x\geq 3\] or else \[[3,\infty)\] depending on the notation used for the interval

OpenStudy (anonymous):

as for the range, the square root symbol means the positive square root, so the range is \[y\geq 0\] or \[[0,\infty)\] again depending on the notation used

OpenStudy (anonymous):

I am sorry I am totally lost. Ithas been several weeks since we covered this. ThankYou for your help I am on problem 29 out of 51 for my final. I am so worried I am not going to pass.... could you help me with the function.... please

OpenStudy (anonymous):

sorry the range of function

OpenStudy (anonymous):

the range is \([0,\infty)\)

OpenStudy (anonymous):

Please can yo help me with another like that one. I just dont get it?